OpenStudy (anonymous):

What is the domain and range for the function F(x) =150(1.02)^x

OpenStudy (anonymous):

ok

OpenStudy (rational):

it seems the graph is staying above x axis ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

which means all real numbers?

OpenStudy (rational):



OpenStudy (rational):

the graph is staying above x axis so the range is all positive numbers (> 0)

OpenStudy (anonymous):

ok

OpenStudy (rational):

domain is all real numbers

OpenStudy (anonymous):

so i was correct

OpenStudy (anonymous):

how would you write that using the −∞≤x≤∞

OpenStudy (anonymous):

thats what i have a hard time doing

OpenStudy (rational):

domain is all real numbers : \(-\infty \lt x\lt \infty \) range is all positive numbers : \(0\lt y \lt \infty\)

OpenStudy (anonymous):

ok thank you so much!!!!!!

OpenStudy (rational):

yw

OpenStudy (anonymous):

Can you help me something else @rational

OpenStudy (rational):

sure tag me

OpenStudy (anonymous):

Using your function f(x) from question 2. Demonstrate and explain how to find the average rate of change between year 3 and 5, and between year 5 and 7. Explain what the average rate of change represents to the frog population. how do you find two points? to find out the ROC

OpenStudy (rational):

you need to provide me that "question 2" also

OpenStudy (anonymous):

F(x) =150(1.02)^x

OpenStudy (rational):

okay plugin x = 3 F(3) = 150(1.02)^3 = ?

OpenStudy (anonymous):

oh never mind i can do it can you check my work for me when i am done?

OpenStudy (rational):

sounds good!

OpenStudy (anonymous):

is it .0214364015??

OpenStudy (anonymous):

@rational

OpenStudy (anonymous):

for year 3 to 5

OpenStudy (anonymous):

and for 5 to 7 is it .0223024323

OpenStudy (rational):

wait

OpenStudy (rational):

for year3 to year5 im getting 3.215 http://www.wolframalpha.com/input/?i=%28150*1.02%5E5-150*1.02%5E3%29%2F2

OpenStudy (rational):

check ur work again

OpenStudy (anonymous):

oh your right thanks

OpenStudy (rational):

np